# Motivic Steenrod operations at the characteristic via infinite ramification

### v1

@@ Appendix C Deglise–Jin–Khan fundamental class in higher Chow groups
Given a regular embedding $i\colon Z\hookrightarrow X$ of codimension $r$, Deglise–Jin–Khan construct a fundamental class $\eta_{i}\in E^{\mathrm{BM}}(Z/X,-\mathcal{N}_{i})$ in the bivariant $E$-theory group of $i$, where $E\in\mathrm{SH}_{X}$. Here, we explicitly compute this class for the motivic cohomology spectrum $H\mathbb{Z}^{\mathbb{A}^{1}}$ when $X$ is smooth over a Dedekind domain $A$. In this situation, the localization theorem of Levine [Lev01, Theorem 1.7] identifies $H\mathbb{Z}^{\mathbb{A}^{1},\mathrm{BM}}(Z/X,-\mathcal{N}_{i})\simeq H\mathbb{Z}^{\mathbb{A}^{1},\mathrm{BM}}(-r)(Z/X)$ with the cycle complex of Bloch $z^{0}(Z,\bullet)\in D(\mathbb{Z})$. The homology groups of this complex are the Higher chow groups of $Z$, $\mathrm{CH}^{r}(Z,n)$, which are isomorphic to the motivic cohomology of $Z$ if $Z$ is either regular and equicharacteristic, or smooth over $A$. We refer to [Gei05] for an introduction to motivic cohomology from the cycle-theoretic point of view.
@@
Figure 1. Illustration of the class $\{t\}\in\mathrm{CH}^{1}(\mathbb{G}_{m},1)=H^{1}_{\mathbb{A}^{1}}(\mathbb{G}_{m},\mathbb{Z}(1))$. The cycles $[A]$ and $[B]$ are the graphs of the two functions $\lambda\to\lambda$ and $\lambda\to\lambda^{-1}$ from $\mathbb{G}_{m}$ to $\mathbb{A}^{1}$, respectively. The two cycles are elements of $z^{1}(\mathbb{G}_{m},1)$ because each of them meets the two faces, represented by the horizontal lines at 00 and $1$, transversely. The difference $[B]-[A]$ is a cycle in the cycle complex, and therefore it gives rise to a class $\{t\}:=[B]-[A]\in\mathrm{CH}^{1}(\mathbb{G}_{m},1)$.
@@
    $\{t\}$ is sent to the cycle $[\mathrm{Spec}(\mathbb{Z})]\in z^{0}(\mathrm{Spec}(\mathbb{Z}),\bullet)$ by the connecting map $\partial^{\prime}$ of the localization sequence
@@
The validity of the second condition is evident, as $B-A$ pulls back to 0 as an algebraic cycle. The validity of the second condition follows from the fact that the algebraic cycle $B-\bar{A}\in z(\mathbb{A}^{1},1)$ intersects the vertical axis at a single point—the origin. Above, $\bar{A}$ is the closure of the algebraic cycle $A\subset\mathbb{G}_{m}\times\mathbb{A}^{1}$ inside $\mathbb{A}^{1}\times\mathbb{A}^{1}$. Thus, by the definition of $\partial^{\prime}$, we have that $\partial^{\prime}\{t\}=[\mathbb{Z}]$ as desired. ∎
@@
By the discussion of the paragraph preceding the statement, it suffices to show that $\sigma_{i}([X])=[\mathcal{N}_{i}]$. To see this, we notice that the closure of the cycle $B\times X-A\times X$ inside $\mathbb{A}^{1}\times D_{i}X$ pulls backs to 0 along the face at $1$, and to $-[\mathcal{N}_{i}]\in z^{1}(D_{i}X;\bullet)$. Thus, by the definition of the connecting map $\partial$ in Eq. 36, we see that $\sigma_{i}([X])=[\mathcal{N}_{i}]$. ∎

### v2

@@ Appendix C Deglise–Jin–Khan fundamental class in higher Chow groups
Given a regular embedding $i\colon Z\hookrightarrow X$ of codimension $r$, Deglise–Jin–Khan [DJK18] construct a fundamental class $\eta_{i}\in E^{\mathrm{BM}}(Z/X,-\mathcal{N}_{i})$ in the bivariant $E$-theory group of $i$, where $E\in\mathrm{SH}_{X}$. Here, we explicitly compute this class for the motivic cohomology spectrum $H\mathbb{Z}^{\mathbb{A}^{1}}$ when $X$ is smooth over a Dedekind domain $A$. In this situation, the localization theorem of Levine [Lev01, Theorem 1.7] identifies $H\mathbb{Z}^{\mathbb{A}^{1},\mathrm{BM}}(Z/X,-\mathcal{N}_{i})\simeq H\mathbb{Z}^{\mathbb{A}^{1},\mathrm{BM}}(-r)(Z/X)$ with the cycle complex of Bloch $z^{0}(Z,\bullet)\in D(\mathbb{Z})$. The homology groups of this complex are the Higher chow groups of $Z$, $\mathrm{CH}^{r}(Z,n)$, which are isomorphic to the motivic cohomology of $Z$ if $Z$ is either regular and equicharacteristic, or smooth over $A$. We refer to [Gei05] for an introduction to motivic cohomology from the cycle-theoretic point of view.
@@
Figure 1. Illustration of the class $\{t\}\in\mathrm{CH}^{1}(\mathbb{G}_{m},1)=H^{1}_{\mathbb{A}^{1}}(\mathbb{G}_{m},\mathbb{Z}(1))$. The cycles $[A]$ and $[B]$ are the graphs of the two functions $\lambda\to\lambda$ and $\lambda\to\lambda^{-1}$ from $\mathbb{G}_{m}$ to $\mathbb{A}^{1}$, respectively. The two cycles are elements of $z^{1}(\mathbb{G}_{m},1)$ because each of them meets the two faces, represented by the horizontal lines at 0 and $1$, transversely. The difference [B][A][B]-[A] is a cycle in the cycle complex, and therefore it gives rise to a class $\{t\}:=[B]-[A]\in\mathrm{CH}^{1}(\mathbb{G}_{m},1)$.
@@
    $\{t\}$ is sent to the cycle $\mathrm{Spec}(\mathbb{Z})]\in z^{0}(\mathrm{Spec}(\mathbb{Z}),\bullet)$ by the connecting map $\partial^{\prime}$ of the localization sequence
@@
The validity of the second condition is evident, as BAB-A pulls back to 0 as an algebraic cycle. The validity of the second condition follows from the fact that the algebraic cycle $B-\bar{A}\in z(\mathbb{A}^{1},1)$ intersects the vertical axis at a single point—the origin. Above, $\bar{A}$ is the closure of the algebraic cycle $A\subset\mathbb{G}_{m}\times\mathbb{A}^{1}$ inside $\mathbb{A}^{1}\times\mathbb{A}^{1}$. Thus, by the definition of $\partial^{\prime}$, we have that $\partial^{\prime}\{t\}=[\mathbb{Z}]$ as desired. ∎
@@
By the discussion of the paragraph preceding the statement, it suffices to show that $\sigma_{i}([X])=[\mathcal{N}_{i}]$. To see this, we notice that the closure of the cycle $B\times X-A\times X$ inside $\mathbb{A}^{1}\times D_{i}X$ pulls backs to 0 along the face at $1$, and to $\mathcal{N}_{i}]\in z^{1}(D_{i}X;\bullet)$. Thus, by the definition of the connecting map $\partial$ in Eq. 38, we see that $\sigma_{i}([X])=[\mathcal{N}_{i}]$. ∎